OpenStudy (anonymous):

Fiona installed a new pool for $9,940 using a 12-month deferred payment plan with an interest rate of 22.53%. What is the balance after the deferment period if payments of $397 are made each month?

OpenStudy (anonymous):

Anyone want to help me out?

OpenStudy (anonymous):

this is the third of these, i am wondering if you are given a formula to use it would make it a lot easier

OpenStudy (anonymous):

lol hey whats up satellite73 =D

OpenStudy (anonymous):

a=p(1+1/n)^nt...............?

jimthompson5910 (jim_thompson5910):

Hint: Use the formula A = P(1 + r/n)^(nt) If the interest is compounded monthly (most likely the case), then n = 12

OpenStudy (anonymous):

yeah but that doesn't really do it

OpenStudy (anonymous):

that tells you what you would get after \(t\) years if you invest \(P\) at a rate of \(r\) compounded \(n\) times per year

jimthompson5910 (jim_thompson5910):

which is how credit card balances are computed (if compound interest is used)

OpenStudy (anonymous):

this is a different question how much is left on the loan

OpenStudy (anonymous):

$7,761.84 $9,940.00 $5,176.00 $12,425.84

jimthompson5910 (jim_thompson5910):

keep in mind no payments are made during the deferred payment period

OpenStudy (anonymous):

but the interest changes, because the principle changes every month

jimthompson5910 (jim_thompson5910):

which is why compound (instead of simple) interest is used

OpenStudy (anonymous):

i know this are the choices but these problems make no sense to me

OpenStudy (anonymous):

these*

OpenStudy (anonymous):

maybe i am confused, but i understand it to mean this: if you pay $397 each month, how much is left on the loan if you are charged no interest, then it is easy if you are charged 22.53% divided by 12 each month, then the principle of the loan changes

jimthompson5910 (jim_thompson5910):

no payments are made during the deferred payment period

OpenStudy (anonymous):

it is not the same as saying "how much would 9,940 be worth at 22.53% in one year?"

jimthompson5910 (jim_thompson5910):

sure you can make payments, but you are allowed to not make them We're assuming no payments are made, which is why the balance goes up

jimthompson5910 (jim_thompson5910):

basically this is saying that you bought an item for some specified amount of money, then you wait 12 months or 1 year til you make payments during that period, the balance goes up because interest is added

OpenStudy (anonymous):

@jim_thompson5910 you may certainly be right, but then why day What is the balance after the deferment period if payments of $397 are made each month?

OpenStudy (anonymous):

*say

jimthompson5910 (jim_thompson5910):

because problems love to throw extra/useless info a lot of the time

OpenStudy (anonymous):

i take this to mean that you pay $397 each month for the first year

jimthompson5910 (jim_thompson5910):

and keep in mind that some other problems like to ask follow up questions ex: when will the balance be fully paid off?

OpenStudy (anonymous):

so how would i set this up then?

jimthompson5910 (jim_thompson5910):

you can, but you don't have to...because the payments are deferred

OpenStudy (anonymous):

again you may be right, but i am sceptical in real life if you have deferred payment you can start buying down the principle right away

jimthompson5910 (jim_thompson5910):

yes, you *can*, but again don't have to you can't get away with not paying down credit debt if it wasn't deferred because you would eventually be legally responsible for doing so (or you would have your credit ruined) with deferred plans, you can wait til the deferred payment period is over when you start making your first payments

jimthompson5910 (jim_thompson5910):

but you're 100% correct, it's smart to pay down your debt asap this is esp true at such a high interest rate

jimthompson5910 (jim_thompson5910):

to start Osniel, use the formula A = P(1 + r/n)^(n*t) where P = principal (amount invested...which in this case is the original balance) r = interest rate n = compounding frequency t = time in years

OpenStudy (anonymous):

i am still confused why the word "if" What is the balance after the deferment period IF payments of $397 are made each month

jimthompson5910 (jim_thompson5910):

but saying "deferment period" implies that payments were deferred

OpenStudy (anonymous):

is it really a red herring?

jimthompson5910 (jim_thompson5910):

why mention there is a deferred period? why not leave that part out

OpenStudy (anonymous):

i thought it meant you made those payment, in which case it would be a totally different formula, but i will believe you

OpenStudy (anonymous):

A=9,940+(0.2253/12)^12*0.2253? is this right

jimthompson5910 (jim_thompson5910):

yes if you made the payments, then that formula wouldn't apply

jimthompson5910 (jim_thompson5910):

more like A = 9940(1 + 0.2253/12)^(12*1)

OpenStudy (anonymous):

oh so is it 12425

jimthompson5910 (jim_thompson5910):

12,425.84 yes

OpenStudy (anonymous):

thank you both of you you two are great teachers

OpenStudy (anonymous):

thanks jim

jimthompson5910 (jim_thompson5910):

np